OpenStudy (anonymous):

Part 1: Factor each trinomial below. Please show your work and check your answer. (1 point each) x2 – 8x + 15 = (x - 5)(x - 3) Check = x^2 + -5x + -3x + 15 = x^2 - 8x + 15 a2 – a – 20 (a - 5)(a + 4) Check = a^2 -5a + 4a -20 = a^2 - a - 20 a2 – 5a – 20 Prime. a2 + 12ab + 27b2 = (a + 2b)( 2a2 + 30a + 100

OpenStudy (anonymous):

I did the first one and the second on if you could check it that would be great and i did the 3 one two and did some of the 4th one...If you could just check them. and i need help with the last one.

OpenStudy (anonymous):

woah, big problem

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

i only need help with the second to last and last ones...

OpenStudy (anonymous):

and if you could check the others...

Parth (parthkohli):

2a2 + 30a + 100 2a2 + 20a + 10a + 100 2a(a + 10) + 10(a + 10) (2a + 10)(a + 10) is the answer

OpenStudy (blockcolder):

a^2+12ab+27b^2 =a^2+9ab+3ab+27b^2 =a(a+9b)+3b(a+9b) =(a+3b)(a+9b)